$SPAD/src/input eqtbl.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool eqtbl.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
e: EqTable(List Integer, Integer) := table()
--R
--R
--R   (1)  table()
--R                                         Type: EqTable(List(Integer),Integer)
--E 1
--S 2 of 6
l1 := [1,2,3]
--R
--R
--R   (2)  [1,2,3]
--R                                                  Type: List(PositiveInteger)
--E 2
--S 3 of 6
l2 := [1,2,3]
--R
--R
--R   (3)  [1,2,3]
--R                                                  Type: List(PositiveInteger)
--E 3
--S 4 of 6
e.l1 := 111
--R
--R
--R   (4)  111
--R                                                        Type: PositiveInteger
--E 4
--S 5 of 6
e.l2 := 222

--R
--R
--R   (5)  222
--R                                                        Type: PositiveInteger
--E 5
--S 6 of 6
e.l1
--R
--R
--R   (6)  111
--R                                                        Type: PositiveInteger
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5